OpenStudy (anonymous):

help? Which of the following is the solution of log x - 50.001 = -3 ? x = 5 x = 7 x = 13 x = 15

OpenStudy (anonymous):

it is \(\log_x(50.001)=-3\) ???

OpenStudy (anonymous):

something goofy about this problem

OpenStudy (anonymous):

If that is the problem, I believe all the choices are wrong

OpenStudy (anonymous):

no sorry the five is with the x

OpenStudy (anonymous):

lets try \[\log_{5x}(.001)=-3\] maybe ?

OpenStudy (anonymous):

its log x-5 together then 0.001=-3

OpenStudy (anonymous):

once more with feeling \[\log_{x-5}(.001)=-3\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

kk so since \(10^{-3}=0.001\) you have \(x-5=10\) and you can solve for \(x\) in one easy step

OpenStudy (anonymous):

can you show me how? I really don't know how to do these..

OpenStudy (anonymous):

this is a tricky one because there is no real algebra of even log way to solve it you have \[\log_{x-5}(0.001)=-3\] which is the same as writing in exponential form \[(x-5)^{-3}=0.001\]

OpenStudy (anonymous):

at this point you just sort of have to know that since \(10^{-3}=\frac{1}{10^3}=\frac{1}{1000}=0.001\) that this means \(x-5\) must be equal to \(10\)

OpenStudy (anonymous):

in other words you were really just supposed to "know" it, you can't really "solve" for it in any case if \(x-5=10\) then \(x=15\)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw